OpenStudy (anonymous):

what is the solution to the absolute value equation? 2[x+2]-1=5

jigglypuff314 (jigglypuff314):

to solve for x: 2|x+2|-1=5 first isolate the absolute value so add 1 to both sides +1 +1 2|x+2| = 6 then divide by 2 on both sides /2 /2 |x+2| = 3 then you can separate both sides into x+2=3 and -(x+2)=3 -x - 2 = 3 then solve for x for both...

OpenStudy (anonymous):

X=1